### v1

# GASP: Gaussian Splatting for Physic-Based Simulations

Abstract
Physical simulation is paramount for modeling and utilization of 3D scenes in various real-world applications. However, its integration with state-of-the-art 3D scene rendering techniques such as Gaussian Splatting (GS) remains challenging. Existing models use additional meshing mechanisms, including triangle or tetrahedron meshing, marching cubes, or cage meshes. As an alternative, we can modify the physically grounded Newtonian dynamics to align with 3D Gaussian components. Current models take the first-order approximation of a deformation map, which locally approximates the dynamics by linear transformations. In contrast, our Gaussian Splatting for Physics-Based Simulations (GASP) model uses such a map (without any modifications) and flat Gaussian distributions, which are parameterized by three points (mesh faces). Subsequently, each 3D point (mesh face node) is treated as a discrete entity within a 3D space. Consequently, the problem of modeling Gaussian components is reduced to working with 3D points. Additionally, the information on mesh faces can be used to incorporate further properties into the physical model, facilitating the use of triangles. Resulting solution can be integrated into any physical engine that can be treated as a black box. As demonstrated in our studies, the proposed model exhibits superior performance on a diverse range of benchmark datasets designed for 3D object rendering. Our project page is at: https://waczjoan.github.io/GASP/.
**footnotetext: These authors contributed equally to this work
1 Introduction
Figure 1: GASP integrates GS with a physical engine to generate realistic simulations. Essentially, GASP utilizes set of triangle meshes to parameterize Gaussian components. This approach reduces modifying Gaussian components to 3D point clouds (consisting of vertices of triangles) processing, enabling efficient and rapid rendering. GASP can model the interaction of objects and work with dynamic scenes.
The recent development of Gaussian Splatting (GS) [13] has greatly influenced the field of computer graphics, enabling the modeling of three-dimensional (3D) scenes from images with annotated camera positions. The GS model represents 3D scenes as a set of 3D Gaussian distributions, which allows for rapid training and rendering. Furthermore, it is relatively straightforward to modify the original GS approach to enable editing [22, 4, 8] and to facilitate the creation of dynamic scenes [21, 11, 14].
Figure 2: GASP produces physical simulations in flat GS. It works with static GS models like GaMeS [22] or dynamic ones like D-MiSo [21]. (In general, GASP can be applied to any GS-based model which uses flat Gaussians.) First, we take the flat Gaussian and convert it to the GaMeS format. Each flat Gaussian component is converted to three points (triangle face). Then, the physical engine is run on a point cloud to obtain trajectories for our 3D model. Finally, we convert points into Gaussian components by reverse GaMeS parametrization.
At the same time, one of the most significant challenges with the GS framework is the incorporation of physics into scenes represented using 3D Gaussian components. It is noteworthy that using a physical engine would pave the way to the implementation of GS in spatial, immersive interfaces such as virtual reality (VR) [12]. Presently, the existing models employ supplementary meshing techniques, including triangle or tetrahedron meshing [2], marching cubes [4], or cage meshes [12]. This approach generates classical mesh-based representations, which are then controlled by a physical engine. Consequently, Gaussian components are adjusted in accordance with mesh-based modifications. While such methods produce satisfactory renderings, they require the implementation of an additional meshing and rendering strategy. Furthermore, modeling tears and cracks represents a challenge, as it is not straightforward to control the behaviors of the mesh when the object is divided. An alternative methodology for addressing the aforementioned issue is to utilize the vanilla GS approach and adapt Newtonian dynamics to accommodate 3D Gaussian components. PhysGaussian [24] employs the first-order approximation of a deformation map. By utilizing such a method, we locally approximate the dynamics through linear transformations. This solution is effective, however, it is contingent upon access to a physical engine in order to apply the modifications and relies on a network approximation.
To address these limitations, we propose the Gaussian Splatting for Physic-Based Simulations (GASP) [^1] The code is avaliable at https://github.com/waczjoan/GASP model (see Fig. 1), which incorporates flat Gaussians [22, 21] with the Material Point Method (MPM) [5] without additional modification (see Fig. 2). Flat Gaussian distributions can be used in GS by applying only minimal changes, such as imposing a zero value on one of the Gaussian components. Such techniques are frequently employed in a multitude of approaches [22, 4, 7], and they facilitate straightforward modifications to the final model. The training and rendering times, as well as the quality of the renderings produced by flat Gaussians, are comparable to those of GS. GASP approach exemplifies the implementation of Newtonian dynamics within the GaMeS framework, thereby illustrating the capacity to integrate these methodologies without the need for external modifications to the physical engine or Gaussian components. More precisely, the model extracts points from GS and applies the Material MPM to control such points. During the simulation process, the Gaussian components associated with each triple of points are recalculated. This process is notably rapid and does not increase rendering costs. At the same time, treating each point as a discrete entity results in the emergence of artifacts. Consequently, we propose a simple modification that utilizes triangles instead of completely disparate points. We use additional rescaling when a triangle undergoes a significant change in size. In practice, GASP can produce high-quality simulations in flat GS, including static and dynamic scenes (see Fig. 3) as well as the interaction of objects (see Fig. 4). This is evidenced by the results of experiments conducted on a range of benchmarks with synthetic (see Fig. 5 ) and real-world data sets (see Fig. 6 ) .
The following constitutes a list of our key contributions:
  1. 1.
    We propose GASP model that incorporates physical properties into a 3D scene representation using GS on both static and dynamic scenes.
  2. 2.
    GASP operates in a direct manner on specific points, thus obviating the necessity for alterations to the physical engine regarded as a black box.
  3. 3.
    GASP does not employ supplementary meshing strategies and instead operates directly with Gaussian distributions.
2 Related Work
The GS representation of 3D scenes is well suited for editing [22, 4, 7] and dynamic scenes modeling [10, 23, 21]. However, the more challenging task is its integration with a physical engine that remains difficult and underexplored.
Figure 3: In this experiment, we added simulated gravity and lowered a rubber teddy bear. As you can see, when the object hits the ground, it changes according to the properties given by the physics engine.
Most of the existing methods use additional strategic tactics to obtain representation well-suited for physical engines. Gaussian Splashing (GSP) [2] is a unified framework that combines 3D Gaussian Splinting (3DGS) and position-based dynamics. During the preprocessing stage, the foreground objects are isolated and reconstructed as a mesh representation. Such a response is then combined with position-based dynamics. Alternatively, we can use cage-based representation [12]. VR-GS [12] combine 3DGS and eXtended Position-based Dynamics (XPBD) [16]. The model starts with scene reconstruction by GS and tetrahedralization. Consequently, we obtain a cage-based representation combined with a physical engine.
All the above models work with an additional meshing strategy. In contrast, PhysGaussian [24] operates directly on Gaussian distributions. The authors of [24] suggest approximating a deformation map using the first-order derivative. This yields a linear transformation capable of modifying Gaussian components. Although this model produces relatively good results, it requires adjustments to the physical engine. In contrast, we propose GASP model which uses a physical engine as a black-box without any modification.
A separate area in the physics modeling involves using diffusion models to generate object dynamics. PhysDreamer [25] introduces a method that provides static 3D objects with interactive dynamics by utilizing object dynamics priors acquired from video generation models. Physics3D [15] learns various physical characteristics of 3D objects using a video diffusion model. DreamPhysics [9] infers physical attributes of 3DGS based on video diffusion priors.
3 GASP Model Overview
Figure 4: GASP reduces the modification of the Gaussian component to 3D point cloud motion. Therefore, we can work with several elements at the same time. In this experiment, we simulate a static mug and a teddy bear made of sand.
Figure 5: We add small elements to large scenes that interact with them. At the top, we add a rubber duck that falls onto the table. At the bottom, we add the sand-based object. Note that GASP is able to model interaction between 3D objects.
Figure 6: In large scenes, we can select part of the object and add physical properties. In this experiment, we simulate the effect of wind to modify the movement of the bonsai tree.
Figure 7: Blender provides a wide range of physical properties and force sources, including gravity and wind, to create realistic and dynamic animations. In this experiment, we present examples of physical animation in object scenes, showing the versatility of animating both small objects, such as a flower on a table, and large objects, such as a fox’s head dominating the scene.
Here, we present a description of the GASP model. It begins with a general GS model, followed by an examination of the modifications made by GaMeS [22] to the original GS model, which facilitates straightforward adaptations. Subsequently, the classical MPM [5] is outlined, and finally, our own approach combining both GS and MPM is detailed.
Gaussian Splatting
The GS technique [13] generates 3D scenes using a collection of three-dimensional Gaussians:
$\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},\Sigma_{i}),\sigma_{i},c_{i})\}_{i=1}^{p},$ (1)
defined by their mean (position) ${\bf m}_{i}$, covariance matrix $\Sigma_{i}$, opacity $\sigma_{i}$, and color $c_{i}$, which is represented using spherical harmonics (SH) [3]. The performance of the GS algorithm is mainly due to its rendering technique, which projects the Gaussian elements into two-dimensional space. Throughout training, all parameters are optimized according to the mean square error (MSE) cost function. Since this training often leads to local minima, GS may utilize additional training methods that include creating, removing, and repositioning components based on the proposed heuristic [13]. This strategy is both fast and efficient.
GaMeS Parametrization of Gaussian Component
In the classical GS model, each element is characterized by a collection of parameters, including a covariance matrix $\Sigma$ which is factored as:
$\Sigma=RSSR^{T},$ (2)
where $R$ is the rotation matrix and $S$ is a diagonal matrix containing the scaling parameters. In GaMeS [22], on the other hand, flat Guassians are used:
$(\mathcal{N}({\bf m},R,S),\sigma,c)$ (3)
where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$. Thus, Gaussian components can be visualized through a triangle-face mesh by utilizing the vertices for their parameterization. We refer to this mapping as $\mathcal{T}(\cdot)$. In practice, this parameterization results in a set of triangles, which is referred to as Triangle Soup [18].
To outline the GaMeS parameterization, consider a single Gaussian component $\mathcal{N}({\bf m},R,S)$. Then its face representation $\mathcal{N}(V)$ is based on a triangle:
$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ (4)
with the vertices defined as:
${\bf v}_{1}={\bf m},\;\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\;\textrm{and}\;\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ (5)
Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component:
$\mathcal{N}(\hat{\bf m},\hat{R},\hat{S})=\mathcal{N}(\mathcal{T}^{-1}(V))$ (6)
through the mean $\hat{\bf m}$, the rotation matrix $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$, and the scaling matrix $\hat{S}=\mathrm{diag}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3})$, where the parameters are defined by the following formulas:
$\hat{\bf m}={\bf v}_{1},\;\;\hat{\bf r}_{1}=\frac{({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})\|},$ (7)
$\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|},\;\;\hat{\bf r}_{3}=\mathrm{orth}({\bf v}_{3}-{\bf v}_{1};{\bf r}_{1},{\bf r}_{2}),$ (8)
$s_{1}=\varepsilon,\;\;\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|,\;\;\textrm{and}\;\;\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle.$ (9)
Here, $\mathrm{orth}(\cdot)$ denotes a single step of the Gram-Schmidt process [1]. Accordingly, the corresponding covariance matrix of a Gaussian distribution is given as:
$\hat{\Sigma}=\hat{R}\hat{S}\hat{S}\hat{R}^{T}.$ (10)
Constitutive Model
In alignment with the methodology delineated by [24], the tenets of continuum mechanics form the foundation of the constitutive model. The fundamental principles of continuum mechanics entail the representation of material deformation through the use of a map $\phi$, which transforms points from the initial undeformed material space $X$ to the subsequently deformed world space $x=\phi(X,t)$ at time $t$. This process yields the following deformation gradient:
$F(X,t)=\nabla_{X}\phi(X,t),$ (11)
which is used to measure local rotation and strain. The deformation function $\phi$ is subject to changes in accordance with the principles of conservation of mass, conservation of momentum, and the elasto-plastic constitutive relation:
$\frac{D\rho}{Dt}=0,\;\;\frac{Dv}{Dt}=\frac{1}{\rho}\nabla\cdot\sigma+f,$ (12)
and
$\sigma=\frac{1}{\det(F)}\frac{\partial\Psi}{\partial F^{E}}({F^{E}})^{T}.$ (13)
In this context, $\rho$ and $v$ denote the density and velocity as functions of time $t$ and position $x$, while $\sigma$ is the Cauchy stress tensor, which includes both pressure and shear, and $\nabla\cdot\sigma$ denotes its divergence. Furthermore, $f$ represents a vector that includes all accelerations caused by body forces (e.g. gravitational acceleration). On the other hand, $\Psi$ is the strain energy density function of $F$, which quantifies the amount of non-rigid deformation. Note that in our work, we use a plasticity model in which the deformation gradient is multiplicatively decomposed (following some yield stress condition) into an elastic part $F^{E}$ and a plastic part $F^{P}$, so that $F=F^{E}F^{P}$.
Figure 8: GASP approach operates directly on Gaussian representations without altering the physics-based deformation network. Initially, the 3D objects are depicted as flat Gaussians. GaMeS then converts these Gaussians into triplets (mesh faces). The simulation engine transforms these points following the MPM model. Subsequently, the new points are reassembled into triangles and then into Gaussians. This process is extremely fast, yielding realistic renderings.
Material Point Method
The fundamental premise of the MPM is the utilization of particles, or material points, for the tracking of mass, momentum, and deformation gradient (here, we track rather $F^{E}$ than $F$). From one perspective, the Lagrangian approach enables the monitoring of time-varying characteristics, such as position, velocity, and deformation gradient, in alignment with the mass conservation principle for each particle. This approach ensures the constancy of the overall mass. Conversely, the lack of mesh connectivity between particles represents a substantial challenge in calculating derivatives, which are crucial for stress-based force evaluation (see Equation (13)) and are, in turn, a fundamental component of the momentum conservation principle. This issue can be addressed by employing a regular Eulerian grid as a background and discretizing the Cauchy stress tensor divergence $\nabla\cdot\sigma$ with the use of interpolation functions over the grid in a manner analogous to the standard Finite Element Method (FEM), utilizing the weak form. In this approach, the grid basis functions are dyadic products of one-dimensional cubic B-splines [19].
The MPM algorithm is a three-stage process comprising particle-to-grid transfer, grid update, and grid-to-particle transfer, which are applied in a loop [24]. In the initial stage, mass and momentum are transferred from the particle to the grid. In the subsequent stage, grid velocities are updated based on forces at the next timestep. In the final stage, the updated grid velocities are transferred back to the particles, and then new quantities of the particles are computed.
GASP – Physics in GaMeS
This subsection presents a description of the GASP model. Since the model employs the GaMeS representation of a 3D scene, we postulate that the input comprises a collection of Gaussians:
$\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},R_{i},S_{i}),\sigma_{i},c_{i})\}_{i=1}^{p},$ (14)
with means ${\bf m}_{i}$, scaling matrices $S_{i}$, and rotation matrices $R_{i}$. Furthermore, we assume the existence of a deformation map $\phi(X,t)$, which models the behavior of 3D objects. In order to obtain $\phi(X,t)$, it is necessary to have a point cloud that describes 3D objects. As our method employs the GaMeS reparametrization of GS, we are able to transform the mean Gaussian parameters and the covariance matrix into a mesh (triangle soup). In such an instance, we may consider the system $(\mathcal{V},\mathcal{E})$, where:
$\mathcal{V}=\{V_{i}=\mathcal{T}({\bf m}_{i},R_{i},S_{i})\}_{i=1}^{p}$ (15)
denotes a collection of vertices, and $\mathcal{E}$ represents the edges obtained via GaMeS. Our physical-based deformation network operates on each point of $V_{i}$ individually. Subsequently, the vertices of the triangle soup can be used as input for the MPM algorithm in conjunction with tools such as the Blender or Taichi libraries (for further details, see the experimental section). It is essential to note that our solution does not entail any modifications to the MPM, in contrast to the Taylor approximation employed in [24]. Consequently, in each time step, the following is obtained:
$\mathcal{V}_{t}=\{\phi(V_{i},t)\colon V_{i}\in\mathcal{V}\}.$ (16)
The mesh in time $t$ is then defined by the collection of vertices and edges $(\mathcal{V}_{t},\mathcal{E}_{t})$. In contrast, the reverse parametrization provided by GaMeS can be employed to obtain Gaussian Splatting:
$\mathcal{G}_{t}=\{(\mathcal{N}(\mathcal{T}^{-1}(\phi(V_{i},t))),\sigma_{i},c_{i})\}_{i=1}^{p}.$ (17)
It is important to note that processing each vertex of a triangle soup individually is a highly rapid process that treats the physical engine as an opaque black box. The schema of such processes is presented in Fig. 8. Our method is compatible with any physics engine that operates on three-dimensional points. However, the processing of all points separately may yield artifacts. When all points are treated independently during simulation, some elements exhibit considerable fluctuations in velocity. This phenomenon leads to the formation of extensive triangles (mesh faces), which in turn give rise to prominent Gaussian components, as illustrated in Fig. 9. To mitigate this effect, our approach employs a straightforward yet highly effective technique (correction algorithm). Specifically, we monitor the distance between vertices of the face representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ of a Gaussian component $\mathcal{N}({\bf m},R,S)$. It should be noticed that in the case of GaMeS transformation, the first vertex is equal to the center of the distribution (${\bf v}_{1}={\bf m}$), while the two subsequent vertices are located on the main components of the ellipses that describe the distribution (${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$). Consequently, if the distance between the transformed mesh nodes is considerably greater than that of the original nodes, i.e.:
$\|\phi({\bf v}_{1},t)-\phi({\bf v}_{i},t)\|>\alpha\|{\bf v}_{1}-{\bf v}_{i}\|\;\;\mbox{for}\;\;i=2,3,$ (18)
where $\alpha$ is a hyperparameter, we automatically assert that the value of $s_{i}$ should be identical to that of the original value times $\alpha$. This is illustrated in Fig. 10. It should be noted that throughout the full simulation, both the mesh and GS are accessible, enabling real-time rendering.
Figure 9: Comparison of physical animation models, illustrating the effects of applying an algorithm for correcting points used in simulation. The left side demonstrates a simulation without the correction algorithm, while the right side shows the enhanced accuracy and stability achieved using Taichi elements with the correction algorithm applied.
Figure 10: GASP works directly on points cloud extracted from flat Gaussians. In practice, physical engine moves all particles independently, which can produce artifacts. Therefore, we use a correction algorithm to control the distance of points from each other. If some points are moved extremely far, we automatically correct such behaviors by clipping parameters $S$ in GS.
GASP for Dynamic Scene
GASP is a natural combination of a physical engine dedicated to points and Gaussian Splatting with flat Gaussians. The reparametrization in GaMeS can be applied to a multitude of disparate methodologies, with a particular focus on flat Gaussian methods [4, 7]. Moreover, it can be employed for dynamic Gaussian Splatting. In [21], D-MiSo model is introduced, which uses flat Gaussians for the modeling of dynamic scenes. Such representations can be utilized in conjunction with the aforementioned approach. In the D-MiSo framework, two distinct types of Gaussians and two deformable networks are employed. In the process of inference, only sub-Gaussians are employed, which are modified by deformable function composition $\psi(v,s)$, i.e.,
$\mathcal{G}_{\text{D-MiSo}}=\left\{\left(\mathcal{N}_{sub}\left(\psi(V_{j}^{i},s)\right),\sigma_{i},c_{i}\right)\right\}_{i=1}^{p},$ (19)
where $s$ represents a time value derived from the dynamics observed in the scenes. In our solution, we can incorporate a deformation map $\phi(X,t)$, which originates from the physics simulator. Consequently, GASP for the dynamic scene is given as follows:
$\mathcal{G}_{t}=\left\{\left(\mathcal{N}_{sub}\left(\phi(\psi(V_{j}^{i},s),t)\right),\sigma_{i},c_{i}\right)\right\}_{i=1}^{p},$ (20)
where $s$ represents time in the dynamic scene and $t$ denotes time in the simulation.
4 Experiments
This experimental section demonstrates that GASP is a universal method that can be applied to both small objects and large scenes. Furthermore, it allows for the modeling of interactions between objects and breaking them into pieces. GASP is also capable of working with dynamic scenes.
Static Objects
The Shiny Dataset [20] contains examples of static objects with varying textures. In such a case, all Gaussian components are converted to GaMeS-based representations, after which training of the physical engine is conducted. Specifically, GASP is employed to move the these representations in accordance with a dynamic function. As can be observed in Fig. 1 and Fig. 3, the model is capable of producing high-quality animation renders. In Tab. 1(a), we compare the FID score obtained by GS and GaMeS on static scenes to GASP during simulations. As we can see, our simulation obtained a slightly worse score but still with high-quality renders.
Figure 11: Controlling animations using physics engines allows for application to dynamic objects. This figure compares animations of a dynamic object at three different times under the same physical parameters. This experiment uses GASP integrated with the D-MiSo model [21].
Dynamic Objects
We show that our pipeline also works on dynamic scenes using sample objects from the D-NeRF dataset [17]. We use D-MiSo [21] to model dynamic scenes by flat Gaussians. Then we select random frame and use physical engine. Fig. 1 and Fig. 11 present how the objects change according to physical engine and timeline. In Tab. 1(b), we present the FID score in three different time-steps of dynamic scenes. As we can see, our renders obtain similar scores, which means that our model produces consistent simulations.s.
Multiple Objects
Our model can model the interaction of static objects. In this case, each element of the scene is trained separately. Also, the properties of each component are specified by different parameters. In Fig. 4 we show a sand-based teddy bear and a glass cup. As we can see, the sand falls smoothly into the cup. Let us take the example of interaction between two objects. In Fig. 9 we show how our model handles the formation of artifacts. In Fig. 5 we integrate small duck objects into a full scene. As we can see, the ducks bounce off the table and fall to the ground. Fig. 1 demonstrates the application of physics to two interacting objects: a falling ball and a car reacting to its impact.
Breaking Objects
GASP uses triangle soup to represent flat Gaussians, so we can easily break objects into pieces. In Fig. 12 we show how a cup can be broken. It is important to note that we need a special physics engine to model such a situation. In this example we use additional information about the mesh to train GaMeS on mesh faces (for details see [22]). In this case, Gaussians lie perfectly on the mesh. The initial Gaussian splatting must be trained with additional conditions, but the final GASP modifications are applied analogously to the previous examples (except that here the model is parameterized by mesh).
Figure 12: In this experiment, we create a simulation of a shattering cup using physics simulations. The GaMeS model is used to create a scenario where Gaussians parameters can be adjusted using the mesh.
Model of Physics
We demonstrate that physics modeling can function as a black box by using two different simulation approaches.
Taichi__\__elements [^2] https://github.com/taichi-dev/taichi__\__elements is a highly efficient physics engine for the simulation of multi-material continuum mechanics, originally developed in Python using taichi[6], which is designed for high-performance parallel numerical computation. It supports multiple materials, including water, elastic objects, snow, and sand.
Blender [^3] https://www.blender.org version 3.6 is the most popular open source application for a variety of 3D computer graphics tasks. Our GaMeS-based representation can be imported into Blender, where it can be seen as a triangle soup. Using tools such as mesh deformation, we were able to smoothly simulate soft body physics. The simulation objects were then exported and rendered using the GS renderer. Due to the high speed of the latter process, this method allows us to render our simulation faster for realistic scenes than in the case of using traditional 3D methods in Blender. The animations created in Blender utilized the ”Lattice” and ”Mesh Deform” modifiers, with the mesh being defined by the user, such as a sphere or a cube.
Table 1: Frechet Inception Distance (FID) is calculated in two settings: static scenes (Bonsai, Fox) and dynamic scene standup. GASP obtains slightly worse FID concerning static model GS and GaMeS; see Tab. 1(a). GASP gives similar FID across dynamic scenes; see Tab. 1(b).
FID ($\downarrow$)
Method Bonsai Fox
3DGS 19.71 51.04
GaMeS 19.64 47.62
Ours 38.33 88.59
(a) Comparison of FID score obtained by classical GS and GaMeS for static scenes and renders on images taken across GASP simulation.
Frame FID ($\downarrow$)
Standup
Start 133.55
Mid 112.79
End 142.90
(b) FID score for three different times obtained by GASP. As we can see, the simulation is consistent across dynamic scenes.
5 Conclusion
In this paper, we introduce GASP, a model that integrates high-quality rendering with accurate physical simulation. By leveraging a GaMeS-based representation of objects, our approach enables the simulation of basic phenomena, such as crushing, and more complex scenarios, including cracking and rainfall. This demonstrates the model’s versatility and effectiveness in handling a wide range of physical interactions within simulated environments. Experimental section, shows that GASP work in many difference scenario and produce high quality simulations.
Limitations
Our model produces artifacts when the simulation changes shapes essentially.

### v3

@@
 Abstract
-Physical simulation is paramount for modeling and utilization of 3D scenes in various real-world applications.
-However, its integration with state-of-the-art 3D scene rendering techniques such as Gaussian Splatting (GS) remains challenging.
+Physics simulation is paramount for modeling and utilizing 3D scenes in various real-world applications.
+However, integrating with state-of-the-art 3D scene rendering techniques such as Gaussian Splatting (GS) remains challenging.
 Existing models use additional meshing mechanisms, including triangle or tetrahedron meshing, marching cubes, or cage meshes.
-As an alternative, we can modify the physically grounded Newtonian dynamics to align with 3D Gaussian components.
+Alternatively, we can modify the physics-grounded Newtonian dynamics to align with 3D Gaussian components.
 Current models take the first-order approximation of a deformation map, which locally approximates the dynamics by linear transformations.
-In contrast, our Gaussian Splatting for Physics-Based Simulations (GASP) model uses such a map (without any modifications) and flat Gaussian distributions, which are parameterized by three points (mesh faces).
-Subsequently, each 3D point (mesh face node) is treated as a discrete entity within a 3D space.
-Consequently, the problem of modeling Gaussian components is reduced to working with 3D points.
-Additionally, the information on mesh faces can be used to incorporate further properties into the physical model, facilitating the use of triangles.
-Resulting solution can be integrated into any physical engine that can be treated as a black box.
-As demonstrated in our studies, the proposed model exhibits superior performance on a diverse range of benchmark datasets designed for 3D object rendering.
-Our project page is at: https://waczjoan.github.io/GASP/.
+In contrast, our GS for Physics-Based Simulations (GASP) pipeline uses parametrized flat Gaussian distributions.
+Consequently, the problem of modeling Gaussian components using the physics engine is reduced to working with 3D points.
+In our work, we present additional rules for manipulating Gaussians, demonstrating how to adapt the pipeline to incorporate meshes, control Gaussian sizes during simulations, and enhance simulation efficiency.
+This is achieved through the Gaussian grouping strategy, which implements hierarchical structuring and enables simulations to be performed exclusively on selected Gaussians.
+The resulting solution can be integrated into any physics engine that can be treated as a black box.
+As demonstrated in our studies, the proposed pipeline exhibits superior performance on a diverse range of benchmark datasets designed for 3D object rendering.
+The project webpage, which includes additional visualizations, can be found at https://waczjoan.github.io/GASP.
 **footnotetext: These authors contributed equally to this work
-1 Introduction
-Figure 1: GASP integrates GS with a physical engine to generate realistic simulations.
-Essentially, GASP utilizes set of triangle meshes to parameterize Gaussian components.
+Figure 1: GASP integrates Gaussian Splatting (GS) with a physics engine to generate realistic simulations.
+Essentially, GASP utilizes a set of triangles to parameterize Gaussian components.
 This approach reduces modifying Gaussian components to 3D point clouds (consisting of vertices of triangles) processing, enabling efficient and rapid rendering.
 GASP can model the interaction of objects and work with dynamic scenes.
+1 Introduction
 The recent development of Gaussian Splatting (GS) [13] has greatly influenced the field of computer graphics, enabling the modeling of three-dimensional (3D) scenes from images with annotated camera positions.
-The GS model represents 3D scenes as a set of 3D Gaussian distributions, which allows for rapid training and rendering.
+The GS model represents 3D scenes as a set of 3D Gaussian distributions, allowing rapid training and rendering.
 Furthermore, it is relatively straightforward to modify the original GS approach to enable editing [5, 9, 22] and to facilitate the creation of dynamic scenes [11, 14, 21].
-Figure 2: GASP produces physical simulations in flat GS.
-It works with static GS models like GaMeS [22] or dynamic ones like D-MiSo [21].
-(In general, GASP can be applied to any GS-based model which uses flat Gaussians.) First, we take the flat Gaussian and convert it to the GaMeS format.
-Each flat Gaussian component is converted to three points (triangle face).
-Then, the physical engine is run on a point cloud to obtain trajectories for our 3D model.
-Finally, we convert points into Gaussian components by reverse GaMeS parametrization.
 At the same time, one of the most significant challenges with the GS framework is the incorporation of physics into scenes represented using 3D Gaussian components.
-It is noteworthy that using a physical engine would pave the way to the implementation of GS in spatial, immersive interfaces such as virtual reality (VR) [12].
-Presently, the existing models employ supplementary meshing techniques, including triangle or tetrahedron meshing [2], marching cubes [4], or cage meshes [12].
-This approach generates classical mesh-based representations, which are then controlled by a physical engine.
-Consequently, Gaussian components are adjusted in accordance with mesh-based modifications.
-While such methods produce satisfactory renderings, they require the implementation of an additional meshing and rendering strategy.
-Furthermore, modeling tears and cracks represents a challenge, as it is not straightforward to control the behaviors of the mesh when the object is divided.
-An alternative methodology for addressing the aforementioned issue is to utilize the vanilla GS approach and adapt Newtonian dynamics to accommodate 3D Gaussian components.
-PhysGaussian [24] employs the first-order approximation of a deformation map.
-By utilizing such a method, we locally approximate the dynamics through linear transformations.
-This solution is effective, however, it is contingent upon access to a physical engine in order to apply the modifications and relies on a network approximation.
-To address these limitations, we propose the Gaussian Splatting for Physic-Based Simulations (GASP) [^1] The code is avaliable at https://github.com/waczjoan/GASP model (see Fig.
-1), which incorporates flat Gaussians [22, 21] with the Material Point Method (MPM) [5] without additional modification (see Fig.
+Existing models employ additional meshing techniques, including triangle or tetrahedron meshing [4], marching cubes [5], or cage meshes [12].
+This approach generates classical mesh-based representations, which are then controlled by a physics engine.
+Consequently, Gaussian components are adjusted following mesh-based modifications.
+Although such methods produce satisfactory renderings, they require the implementation of an additional meshing and rendering strategy.
+Extracting meshes from 3DGS remains a subject of active research, with promising approaches such as SuGaR or 2DGS [8].
+However, obtaining smooth, high-quality meshes for arbitrary objects or scenes is still a challenging task, and artifacts affecting the simulation, such as mesh gaps, may occur.
+Alternatively, PhysGaussian [24] utilizes the vanilla GS approach and adapts Newtonian dynamics to accommodate 3D Gaussian components.
+PhysGaussian employs the first-order approximation of a deformation map.
+Using such a method, it can locally approximate the dynamics through linear transformations.
+This solution is effective, however, it is contingent upon access to a physics engine in order to apply the modifications and relies on a network approximation.
+Due to that, this method is not suitable for use in other popular 3D graphics software such as Blender.
+To address these limitations, we propose the Gaussian Splatting for Physics-Based Simulations (GASP) pipeline (see Fig.
+1), which incorporates flat Gaussians [21, 22] with black box physics engine without special modifications to GS-based model which makes it more flexible compared to other methods.
+In this work, we present simulations created using the Genesis[2], Blender software and the Taichi Elements library (see Fig.
 2).
-Flat Gaussian distributions can be used in GS by applying only minimal changes, such as imposing a zero value on one of the Gaussian components.
-Such techniques are frequently employed in a multitude of approaches [22, 4, 7], and they facilitate straightforward modifications to the final model.
+Additionally, to the best of our knowledge, this is the first demonstration of the integration of a GS-based model with the Genesis engine.
+We also demonstrate that straightforward enhancements, such as implementing the Gaussian hierarchy, can seamlessly integrate into our method, significantly accelerating the simulation process.
+Figure 2: The GASP pipeline is versatile and compatible with various engines, requiring no significant modifications for integration.
+We demonstrate its capabilities with simulations performed using the Genesis engine, Blender software, and Taichi Elements.
+Flat Gaussian distributions can be smoothly employed in GS by modifying Gaussian’s scaling component, so it has a zero value on one axis.
+This representation of GS enables straightforward parametrization and can be easily converted into faces compatible with traditional 3D graphic software.
+Such techniques have been previously used in several approaches [5, 8, 22].
 The training and rendering times, as well as the quality of the renderings produced by flat Gaussians, are comparable to those of GS.
-GASP approach exemplifies the implementation of Newtonian dynamics within the GaMeS framework, thereby illustrating the capacity to integrate these methodologies without the need for external modifications to the physical engine or Gaussian components.
-More precisely, the model extracts points from GS and applies the Material MPM to control such points.
+GASP approach exemplifies the implementation of Newtonian dynamics within the GaMeS framework, thereby illustrating the capacity to integrate these methodologies without needing external modifications to the physics engine or Gaussian components.
+More precisely, the model extracts points from GS and applies the Material Point Method to control such points.
 During the simulation process, the Gaussian components associated with each triple of points are recalculated.
@@
 We use additional rescaling when a triangle undergoes a significant change in size.
-In practice, GASP can produce high-quality simulations in flat GS, including static and dynamic scenes (see Fig.
-3) as well as the interaction of objects (see Fig.
-4).
-This is evidenced by the results of experiments conducted on a range of benchmarks with synthetic (see Fig.
-5 ) and real-world data sets (see Fig.
-6 ) .
+In practice, GASP can produce high-quality simulations in flat GS, including static and dynamic scenes as well as the interaction of objects, see Fig.
+1.
+This is evidenced by the results of experiments conducted on a range of benchmarks with synthetic and real-world data sets.
 The following constitutes a list of our key contributions:
   1. 1.
-    We propose GASP model that incorporates physical properties into a 3D scene representation using GS on both static and dynamic scenes.
+    We propose GASP pipeline that incorporates physical properties into a 3D scene representation using GS-based models on both static and dynamic scenes.
   2. 2.
-    GASP operates in a direct manner on specific points, thus obviating the necessity for alterations to the physical engine regarded as a black box.
+    GASP operates directly on specific points, thus obviating the necessity for alterations to the physics engine regarded as a black box.
   3. 3.
-    GASP does not employ supplementary meshing strategies and instead operates directly with Gaussian distributions.
+    GASP does not employ additional meshing strategies.
+We demonstrate that applying additional Gaussian control rules effectively eliminates artifacts, while leveraging hierarchy significantly accelerates simulations.
 2 Related Work
-The GS representation of 3D scenes is well suited for editing [22, 4, 7] and dynamic scenes modeling [10, 23, 21].
-However, the more challenging task is its integration with a physical engine that remains difficult and underexplored.
-Figure 3: In this experiment, we added simulated gravity and lowered a rubber teddy bear.
-As you can see, when the object hits the ground, it changes according to the properties given by the physics engine.
-Most of the existing methods use additional strategic tactics to obtain representation well-suited for physical engines.
-Gaussian Splashing (GSP) [2] is a unified framework that combines 3D Gaussian Splinting (3DGS) and position-based dynamics.
-During the preprocessing stage, the foreground objects are isolated and reconstructed as a mesh representation.
+The GS representation of 3D scenes is suitable for editing [5, 8, 22] and modeling dynamic scenes [11, 23, 21].
+However, the more challenging task is its integration with a physics engine that remains difficult and underexplored.
+Figure 3: GASP produces physical simulations using flat GS representation.
+It works with static GS models like GaMeS [22] or dynamic ones like D-MiSo [21].
+First, each flat Gaussian component is converted to three points (a triangle face).
+Then, the physics engine is run on a selected point cloud to obtain trajectories for our 3D model.
+Finally, we convert points into Gaussian components by reverse GaMeS parametrization.
+Notably, we also provided some additional rules that can be applied to control Gaussians during simulations.
+Most existing methods use additional strategic tactics to obtain representations better aligned with physics engines’ requirements.
+Gaussian Splashing (GSP) [4] is a unified framework that combines 3D Gaussian Splatting (3DGS) and position-based dynamics.
+During the preprocessing, the foreground objects are isolated and reconstructed as a mesh.
 Such a response is then combined with position-based dynamics.
+Case-specific methods, such as Embodied Gaussians [1], introduce a Gaussian Particle that represents an object through particles in order to maintain its geometric aspects for robotics applications.
 Alternatively, we can use cage-based representation [12].
-VR-GS [12] combine 3DGS and eXtended Position-based Dynamics (XPBD) [16].
+Other methods, such as LITA-GS [26], extract physical attributes from scenes in order to adapt the lighting conditions.
+VR-GS [12] combines 3DGS and eXtended Position-based Dynamics (XPBD) [16].
 The model starts with scene reconstruction by GS and tetrahedralization.
-Consequently, we obtain a cage-based representation combined with a physical engine.
-All the above models work with an additional meshing strategy.
+Consequently, we obtain a cage-based representation merged with a physics engine.
+All of the above models work with an additional meshing strategy.
 In contrast, PhysGaussian [24] operates directly on Gaussian distributions.
@@
 This yields a linear transformation capable of modifying Gaussian components.
-Although this model produces relatively good results, it requires adjustments to the physical engine.
-In contrast, we propose GASP model which uses a physical engine as a black-box without any modification.
-A separate area in the physics modeling involves using diffusion models to generate object dynamics.
+Although this model produces relatively good results, it requires adjustments to the physics engine.
+In contrast, we propose GASP model, which uses a physics engine as a black box without any modification.
+A separate area in physics modeling involves using diffusion models to generate object dynamics.
 PhysDreamer [25] introduces a method that provides static 3D objects with interactive dynamics by utilizing object dynamics priors acquired from video generation models.
 Physics3D [15] learns various physical characteristics of 3D objects using a video diffusion model.
-DreamPhysics [9] infers physical attributes of 3DGS based on video diffusion priors.
-3 GASP Model Overview
-Figure 4: GASP reduces the modification of the Gaussian component to 3D point cloud motion.
-Therefore, we can work with several elements at the same time.
-In this experiment, we simulate a static mug and a teddy bear made of sand.
-Figure 5: We add small elements to large scenes that interact with them.
-At the top, we add a rubber duck that falls onto the table.
-At the bottom, we add the sand-based object.
-Note that GASP is able to model interaction between 3D objects.
-Figure 6: In large scenes, we can select part of the object and add physical properties. In this experiment, we simulate the effect of wind to modify the movement of the bonsai tree.
-Figure 7: Blender provides a wide range of physical properties and force sources, including gravity and wind, to create realistic and dynamic animations.
-In this experiment, we present examples of physical animation in object scenes, showing the versatility of animating both small objects, such as a flower on a table, and large objects, such as a fox’s head dominating the scene.
-Here, we present a description of the GASP model.
+DreamPhysics [10] infers the physical attributes of 3DGS based on video diffusion priors.
+3 GASP pipeline Overview
+Here, we present a description of the GASP pipeline (see Fig.
+3).
 It begins with a general GS model, followed by an examination of the modifications made by GaMeS [22] to the original GS model, which facilitates straightforward adaptations.
 Subsequently, the classical MPM [6] is outlined, and finally, our own approach combining both GS and MPM is detailed.
-Gaussian Splatting
-The GS technique [13] generates 3D scenes using a collection of three-dimensional Gaussians:
-$\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},\Sigma_{i}),\sigma_{i},c_{i})\}_{i=1}^{p},$ (1)
-defined by their mean (position) ${\bf m}_{i}$, covariance matrix $\Sigma_{i}$, opacity $\sigma_{i}$, and color $c_{i}$, which is represented using spherical harmonics (SH) [3].
-The performance of the GS algorithm is mainly due to its rendering technique, which projects the Gaussian elements into two-dimensional space.
-Throughout training, all parameters are optimized according to the mean square error (MSE) cost function.
-Since this training often leads to local minima, GS may utilize additional training methods that include creating, removing, and repositioning components based on the proposed heuristic [13].
-This strategy is both fast and efficient.
+In order to improve visual and time quality, we have also proposed two strategies.
+One is based on the control of the size of Gaussians, and the other is on making appropriate Core Gaussians to create hierarchies (see Fig.
+4).
 GaMeS Parametrization of Gaussian Component
-In the classical GS model, each element is characterized by a collection of parameters, including a covariance matrix $\Sigma$ which is factored as:
-$\Sigma=RSSR^{T},$ (2)
-where $R$ is the rotation matrix and $S$ is a diagonal matrix containing the scaling parameters. In GaMeS [22], on the other hand, flat Guassians are used:
-$(\mathcal{N}({\bf m},R,S),\sigma,c)$ (3)
-where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$.
+Figure 4: In the GASP pipeline, the hierarchical organization of Gaussians through clustering methods allows simulations to focus only on Gaussians at the highest levels of the hierarchy.
+This approach speeds up computations, especially when using particle-based engines such as Taichi.
+In the classical GS model, each element is characterized by a collection of parameters, including a covariance matrix $\Sigma$ which is factored as $\Sigma=RSS^{T}R^{T}$ where $R$ is the rotation matrix and $S$ is a diagonal matrix containing the scaling parameters.
+In GaMeS [22], on the other hand, flat Guassians are used $(\mathcal{N}({\bf m},R,S),\sigma,c),$ where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$.
 Thus, Gaussian components can be visualized through a triangle-face mesh by utilizing the vertices for their parameterization.
@@
 In practice, this parameterization results in a set of triangles, which is referred to as triangle soup [18].
-To outline the GaMeS parameterization, consider a single Gaussian component $\mathcal{N}({\bf m},R,S)$. Then its face representation $\mathcal{N}(V)$ is based on a triangle:
-$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ (4)
-with the vertices defined as:
-${\bf v}_{1}={\bf m},\;\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\;\textrm{and}\;\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ (5)
-Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component:
+To outline the GaMeS parameterization, consider a single Gaussian component $\mathcal{N}({\bf m},R,S)$.
+Then its face representation $\mathcal{N}(V)$ is based on a triangle $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ with the vertices defined as ${\bf v}_{1}={\bf m},\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\textrm{and}\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component:
 $\mathcal{N}(\hat{\bf m},\hat{R},\hat{S})=\mathcal{N}(\mathcal{T}^{-1}(V))$ (1)
@@
 $s_{1}=\varepsilon,\;\;\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|,\;\;\textrm{and}\;\;\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle.$ (4)
-Here, $\mathrm{orth}(\cdot)$ denotes a single step of the Gram-Schmidt process [1]. Accordingly, the corresponding covariance matrix of a Gaussian distribution is given as:
-$\hat{\Sigma}=\hat{R}\hat{S}\hat{S}\hat{R}^{T}.$ (10)
+Here, $\mathrm{orth}(\cdot)$ denotes a single step of the Gram-Schmidt process [3].
+Accordingly, the corresponding covariance matrix of a Gaussian distribution is given as $\hat{\Sigma}=\hat{R}\hat{S}\hat{S}^{T}\hat{R}^{T}.$
+In addition, GaMeS supports distributing Gaussians on meshes when a mesh structure is provided.
+In this setting, Gaussian means are assigned to each mesh face via learnable barycentric coordinates of its vertices, ensuring they remain bound to the mesh under deformations.
+The orientation of each Gaussian is defined by the face normal and two in-plane directions derived from the vertices, allowing consistent updates when the mesh position changes.
+In this work, we focus on the mesh-free case, though the proposed formulations are directly applicable to Gaussians placed on mesh faces.
+$t_{0}$ $t_{i}$ $t_{j}$ $t_{T}$
+2DGS Refer to caption Refer to caption Refer to caption Refer to caption
+GaMeS Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 5: Visual comparison of gravity simulations with sand-like material represented by points.
+First row: direct application of physical simulators to Gaussians obtained from 2DGS, which produces noticeable artifacts.
+Second row: application of physical simulators on triangle faces obtained using the GaMeS parametrization, the pointed artifacts and holes are not as apparent.
 Constitutive Model
@@
 $F(X,t)=\nabla_{X}\phi(X,t),$ (5)
-which is used to measure local rotation and strain.
+, which is used to measure local rotation and strain.
 The deformation function $\phi$ is subject to changes in accordance with the principles of conservation of mass, conservation of momentum, and the elasto-plastic constitutive relation:
@@
 Note that in our work, we use a plasticity model in which the deformation gradient is multiplicatively decomposed (following some yield stress condition) into an elastic part $F^{E}$ and a plastic part $F^{P}$, so that $F=F^{E}F^{P}$.
-Figure 8: GASP approach operates directly on Gaussian representations without altering the physics-based deformation network.
-Initially, the 3D objects are depicted as flat Gaussians.
-GaMeS then converts these Gaussians into triplets (mesh faces).
-The simulation engine transforms these points following the MPM model.
-Subsequently, the new points are reassembled into triangles and then into Gaussians.
-This process is extremely fast, yielding realistic renderings.
 Material Point Method
@@
 In the final stage, the updated grid velocities are transferred back to the particles, and then new quantities of the particles are computed.
+Figure 6: GASP operates directly on the Gaussian triangle representation without altering the physics-based simulation engine.
+Initially, the 3D objects are depicted as flat Gaussians.
+GaMeS then converts these Gaussians into triplets (mesh faces).
+The simulation engine transforms these points following the MPM model.
+Subsequently, the new points are reassembled into triangles and then into Gaussians.
+This process is speedy, yielding realistic renderings.
 GASP – Physics in GaMeS
-This subsection presents a description of the GASP model. Since the model employs the GaMeS representation of a 3D scene, we postulate that the input comprises a collection of Gaussians:
+This subsection presents a description of the GASP pipeline. Since the model employs the GaMeS representation of a 3D scene, we postulate that the input comprises a collection of Gaussians:
 $\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},R_{i},S_{i}),\sigma_{i},c_{i})\}_{i=1}^{p},$ (8)
@@
 Furthermore, we assume the existence of a deformation map $\phi(X,t)$, which models the behavior of 3D objects.
-In order to obtain $\phi(X,t)$, it is necessary to have a point cloud that describes 3D objects.
-As our method employs the GaMeS reparametrization of GS, we are able to transform the mean Gaussian parameters and the covariance matrix into a mesh (triangle soup).
+To obtain $\phi(X,t)$, it is necessary to have a point cloud that describes 3D objects.
+As our method employs the GaMeS reparametrization of GS, we can transform the mean Gaussian parameters and the covariance matrix into a mesh (triangle soup).
 In such an instance, we may consider the system $(\mathcal{V},\mathcal{E})$, where:
@@
 denotes a collection of vertices, and $\mathcal{E}$ represents the edges obtained via GaMeS.
-Our physical-based deformation network operates on each point of $V_{i}$ individually.
-Subsequently, the vertices of the triangle soup can be used as input for the MPM algorithm in conjunction with tools such as the Blender or Taichi libraries (for further details, see the experimental section).
+Our physics-based simulation engine operates at each point of $V_{i}$ individually.
+Subsequently, the vertices of the triangle soup can be used as input for the MPM algorithm in conjunction with tools such as the Blender software, Taichi_elements, or Genesis libraries (for further details, see the experimental section), see Fig.
+2.
 It is essential to note that our solution does not entail any modifications to the MPM, in contrast to the Taylor approximation employed in [24].
@@
 $\mathcal{V}_{t}=\{\phi(V_{i},t)\colon V_{i}\in\mathcal{V}\}.$ (10)
-The mesh in time $t$ is then defined by the collection of vertices and edges $(\mathcal{V}_{t},\mathcal{E}_{t})$.
+The triangle soup in time $t$ is then defined by the collection of vertices and edges $(\mathcal{V}_{t},\mathcal{E}_{t})$.
 In contrast, the reverse parametrization provided by GaMeS can be employed to obtain Gaussian Splatting:
 $\mathcal{G}_{t}=\{(\mathcal{N}(\mathcal{T}^{-1}(\phi(V_{i},t))),\sigma_{i},c_{i})\}_{i=1}^{p}.$ (11)
-It is important to note that processing each vertex of a triangle soup individually is a highly rapid process that treats the physical engine as an opaque black box.
+Our GASP pipeline represents Gaussian components through triangle faces and integrates them with a point-based physics engine.
+Flat Gaussian components are first converted into triangle faces, and the simulation is then performed on their vertices.
+Direct application of standard physics simulators often produces noticeable artifacts (see Fig.
+5).
+In contrast, our approach effectively supports the learned flat Gaussians by transforming them into triangles, ensuring stable simulations.
+Unlike alternative flat Gaussian representations, such as 2DGS, the GaMeS parametrization naturally fits this setting, since each flat Gaussian is inherently represented by a triangle.
+Figure 7: Comparison of physical animation models, illustrating the effects of applying an algorithm to correct points used in simulation.
+The left side demonstrates a simulation without the correction algorithm, while the right side shows the enhanced accuracy and stability achieved using Taichi elements with the correction algorithm applied.
+Figure 8: GASP works directly on points cloud extracted from flat Gaussians.
+In practice, physics engine like Taichi_elements moves all particles independently, which can produce artifacts.
+Therefore, we use a correction algorithm to control the distance of points from each other.
+If some points are moved extremely far, we automatically correct such behaviors by clipping parameters $S$ in GS.
+Gaussian Control Rules
+It is important to note that processing each vertex of a triangle soup individually is a highly rapid process that treats the physics engine as a black box.
 The schema of such processes is presented in Fig.
-8.
+6.
 Our method is compatible with any physics engine that operates on three-dimensional points.
@@
 This phenomenon leads to the formation of extensive triangles (mesh faces), which in turn give rise to prominent Gaussian components, as illustrated in Fig.
-9.
+7.
 To mitigate this effect, our approach employs a straightforward yet highly effective technique (correction algorithm).
 Specifically, we monitor the distance between vertices of the face representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ of a Gaussian component $\mathcal{N}({\bf m},R,S)$.
-It should be noticed that in the case of GaMeS transformation, the first vertex is equal to the center of the distribution (${\bf v}_{1}={\bf m}$), while the two subsequent vertices are located on the main components of the ellipses that describe the distribution (${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$).
+It should be noticed that, in the case of GaMeS transformation, the first vertex is equal to the center of the distribution (${\bf v}_{1}={\bf m}$), while the two subsequent vertices are located on the main components of the ellipses that describe the distribution (${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$).
 Consequently, if the distance between the transformed mesh nodes is considerably greater than that of the original nodes, i.e.:
 $\|\phi({\bf v}_{1},t)-\phi({\bf v}_{i},t)\|>\alpha\|{\bf v}_{1}-{\bf v}_{i}\|\;\;\mbox{for}\;\;i=2,3,$ (12)
-where $\alpha$ is a hyperparameter, we automatically assert that the value of $s_{i}$ should be identical to that of the original value times $\alpha$.
-This is illustrated in Fig.
-10.
-It should be noted that throughout the full simulation, both the mesh and GS are accessible, enabling real-time rendering.
-Figure 9: Comparison of physical animation models, illustrating the effects of applying an algorithm for correcting points used in simulation.
-The left side demonstrates a simulation without the correction algorithm, while the right side shows the enhanced accuracy and stability achieved using Taichi elements with the correction algorithm applied.
-Figure 10: GASP works directly on points cloud extracted from flat Gaussians.
-In practice, physical engine moves all particles independently, which can produce artifacts.
-Therefore, we use a correction algorithm to control the distance of points from each other.
-If some points are moved extremely far, we automatically correct such behaviors by clipping parameters $S$ in GS.
+where $\alpha$ is a hyperparameter, we automatically assert that the value of $s_{i}$ should be identical to that of the original value times $\alpha$. This is illustrated in Fig. 8.
+Gaussians Hierarchy in Simulations
+Figure 9: Blender provides a wide range of physical properties and force sources, including gravity and wind, to create realistic and dynamic animations.
+In this experiment, we present examples of physics animation in object scenes, showing the versatility of animating both small objects, such as a flower on a table, and large objects, such as a fox’s head dominating the scene.
+Achieving high-quality renders with Gaussian Splatting methods typically requires a substantial number of Gaussians.
+Previously, animations were executed directly on Gaussians [24] or using estimated meshes [5].
+However, these simulations demand significant time and computational resources.
+In the Genesis engine [2], this process is optimized by employing simulations facilitated through particles selected via Sampler component, lowering the number of points.
+To achieve a similar mechanism, we employ BIRCH clustering to construct Gaussians called Core-Gaussians for each cluster.
+This enables the creation of a hierarchical structure for Gaussians.
+The simulation is then performed only on Core Gaussians, and then their move pattern is transferred on Sub Gaussians inspired by [21], see Fig.
+4.
+We demonstrate that physics engines can be applied to selected Gaussians without compromising visual quality (see Fig.
+4).
+This approach significantly accelerates the simulation process while maintaining rendering quality.
 GASP for Dynamic Scene
-GASP is a natural combination of a physical engine dedicated to points and Gaussian Splatting with flat Gaussians.
-The reparametrization in GaMeS can be applied to a multitude of disparate methodologies, with a particular focus on flat Gaussian methods [4, 7].
+GASP is a natural combination of a physics engine dedicated to points and Gaussian Splatting with flat Gaussians.
+The reparametrization in GaMeS can be applied to a multitude of disparate methodologies, with a particular focus on flat Gaussian methods.
 Moreover, it can be employed for dynamic Gaussian Splatting.
-In [21], D-MiSo model is introduced, which uses flat Gaussians for the modeling of dynamic scenes.
+In [21], the D-MiSo model is introduced, which uses flat Gaussians for the modeling of dynamic scenes.
 Such representations can be utilized in conjunction with the aforementioned approach.
@@
 where $s$ represents a time value derived from the dynamics observed in the scenes.
-In our solution, we can incorporate a deformation map $\phi(X,t)$, which originates from the physics simulator.
+In our solution, we can incorporate a deformation map $\phi(X,t)$, which comes from the physics simulator.
 Consequently, GASP for the dynamic scene is given as follows:
@@
 4 Experiments
-This experimental section demonstrates that GASP is a universal method that can be applied to both small objects and large scenes.
+This experimental section demonstrates that GASP is a universal method that can be applied to small objects and large scenes alike.
 Furthermore, it allows for the modeling of interactions between objects and breaking them into pieces.
 GASP is also capable of working with dynamic scenes.
+$t_{0}$ $t_{i}$ $t_{T}$
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 10: Jelly cube falling simulation created in Blender using soft-body and cloth modifiers to mimic gel-like behavior.
+In this setup, each cube was trained independently with GaMeS, allowing every cube in the scene to exhibit distinct properties and more realistic collision interactions.
+$t_{0}$ $t_{i}$ $t_{T}$
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 11: Flag movement simulation created in Blender using the Lattice modifier, based on a triangle soup.
 Static Objects
 The Shiny Dataset [20] contains examples of static objects with varying textures.
-In such a case, all Gaussian components are converted to GaMeS-based representations, after which training of the physical engine is conducted.
-Specifically, GASP is employed to move the these representations in accordance with a dynamic function.
+In such a case, all Gaussian components are converted to GaMeS-based representations, after which the physics engine is used to modify their positions.
+Specifically, GASP is employed to move these representations in accordance with a dynamic function.
 As can be observed in Fig.
-1 and Fig.
-3, the model is capable of producing high-quality animation renders.
-In Tab.
-1(a), we compare the FID score obtained by GS and GaMeS on static scenes to GASP during simulations.
-As we can see, our simulation obtained a slightly worse score but still with high-quality renders.
+1, Fig.
+2, Fig.
+9, Fig.
+10, and Fig.
+11 the model can produce high-quality animation renders.
 Figure 12: Controlling animations using physics engines allows for application to dynamic objects.
-This figure compares animations of a dynamic object at three different times under the same physical parameters.
+This figure compares animations of a dynamic object at two different times under the same physical parameters.
 This experiment uses GASP integrated with the D-MiSo model [21].
 Dynamic Objects
-We show that our pipeline also works on dynamic scenes using sample objects from the D-NeRF dataset [17].
-We use D-MiSo [21] to model dynamic scenes by flat Gaussians.
-Then we select random frame and use physical engine.
+We show that our pipeline also works with dynamic scenes using sample objects from the D-NeRF dataset [17].
+We use D-MiSo [21] to model dynamic scenes using flat Gaussians.
+Then, we select a frame and use a physics engine.
 Fig.
 1 and Fig.
-11 present how the objects change according to physical engine and timeline.
+12 present how the objects change according to the physics engine and timeline.
+Multiple Objects
+We show that GASP can model the interaction of static objects, as shown in Fig.
+6.
+In this case, each element of the scene is trained separately.
+In addition, the properties of each component are specified by different parameters.
+Fig.
+1 shows the application of physics to two interacting objects: a falling ball and a car reacting to its impact.
+In Fig.
+2, we integrate small duck objects into a full scene.
+As can be seen, the ducks bounce off the table and fall to the ground.
+In Fig.
+7, we show a sand-based teddy bear and a glass cup.
+As we can see, the sand falls smoothly into the cup.
+In the same figure, we also show how our model handles the formation of artifacts.
+Model of Physics
+We demonstrate that physics modeling can function as a black box by using three different simulation approaches.
+Genesis [^1] https://github.com/Genesis-Embodied-AI/Genesis [2] is a versatile physics simulation platform designed for Robotics, Embodied AI, and Physical AI applications, featuring a robust physics engine, ultra-fast robotics simulation.
+It supports diverse physics solvers including, but not limited to, MPM, FEM, PBD, as well as various material models.
+In our work, we focus on the MPM solver.
+Taichi_\_elements [^2] https://github.com/taichi-dev/taichi_\_elements is a highly efficient physics engine for the simulation of multi-material continuum mechanics, originally developed in Python using taichi [7], which is designed for high-performance parallel numerical computation.
+It supports multiple materials, including water, elastic objects, snow, and sand.
+We also carried out an experiment to visually and numerically compare our method with PhysGaussian (see Fig.
+13).
+Specifically, we trained a GaMeS model and integrated it directly into their engine.
+Their approach performs MPM simulations on Gaussian centers followed by covariance calculations.
+In contrast, we utilized their engine but performed simulations on designated points from Gaussians without additional covariance computations.
 In Tab.
-1(b), we present the FID score in three different time-steps of dynamic scenes.
-As we can see, our renders obtain similar scores, which means that our model produces consistent simulations.s.
-Multiple Objects
-Our model can model the interaction of static objects.
-In this case, each element of the scene is trained separately.
-Also, the properties of each component are specified by different parameters.
-In Fig.
-4 we show a sand-based teddy bear and a glass cup.
-As we can see, the sand falls smoothly into the cup.
-Let us take the example of interaction between two objects.
-In Fig.
-9 we show how our model handles the formation of artifacts.
-In Fig.
-5 we integrate small duck objects into a full scene.
-As we can see, the ducks bounce off the table and fall to the ground.
-Fig.
-1 demonstrates the application of physics to two interacting objects: a falling ball and a car reacting to its impact.
-Breaking Objects
-GASP uses triangle soup to represent flat Gaussians, so we can easily break objects into pieces.
-In Fig.
-12 we show how a cup can be broken.
-It is important to note that we need a special physics engine to model such a situation.
-In this example we use additional information about the mesh to train GaMeS on mesh faces (for details see [22]).
-In this case, Gaussians lie perfectly on the mesh.
-The initial Gaussian splatting must be trained with additional conditions, but the final GASP modifications are applied analogously to the previous examples (except that here the model is parameterized by mesh).
-Figure 12: In this experiment, we create a simulation of a shattering cup using physics simulations.
-The GaMeS model is used to create a scenario where Gaussians parameters can be adjusted using the mesh.
-Model of Physics
-We demonstrate that physics modeling can function as a black box by using two different simulation approaches.
-Taichi__\__elements [^2] https://github.com/taichi-dev/taichi__\__elements is a highly efficient physics engine for the simulation of multi-material continuum mechanics, originally developed in Python using taichi[6], which is designed for high-performance parallel numerical computation.
-It supports multiple materials, including water, elastic objects, snow, and sand.
-Blender [^3] https://www.blender.org version 3.6 is the most popular open source application for a variety of 3D computer graphics tasks.
-Our GaMeS-based representation can be imported into Blender, where it can be seen as a triangle soup.
-Using tools such as mesh deformation, we were able to smoothly simulate soft body physics.
-The simulation objects were then exported and rendered using the GS renderer.
-Due to the high speed of the latter process, this method allows us to render our simulation faster for realistic scenes than in the case of using traditional 3D methods in Blender.
-The animations created in Blender utilized the ”Lattice” and ”Mesh Deform” modifiers, with the mesh being defined by the user, such as a sphere or a cube.
-Table 1: Frechet Inception Distance (FID) is calculated in two settings: static scenes (Bonsai, Fox) and dynamic scene standup.
-GASP obtains slightly worse FID concerning static model GS and GaMeS; see Tab.
-1(a).
-GASP gives similar FID across dynamic scenes; see Tab.
-1(b).
-FID ($\downarrow$)
-Method Bonsai Fox
-3DGS 19.71 51.04
-GaMeS 19.64 47.62
-Ours 38.33 88.59
-(a) Comparison of FID score obtained by classical GS and GaMeS for static scenes and renders on images taken across GASP simulation.
-Frame FID ($\downarrow$)
-Standup
-Start 133.55
-Mid 112.79
-End 142.90
-(b) FID score for three different times obtained by GASP. As we can see, the simulation is consistent across dynamic scenes.
+1 shows that, qualitatively, the FID remains consistent with PhysGaussian.
+Since we use the same engine and the same number of Gaussians, the result is expected and maintains fairness in comparison.
+Tab.
+2 shows that incorporating hierarchical structuring significantly accelerates the simulation process.
+Moreover, visual comparison is included in Fig.
+13.
+Method Vasedeck Fox Teddybear Ficus Bottle
+PhysGaussian 68.66 167.16 227.91 86.63 261.48
+GASP (our) 68.22 168.55 244.39 83.98 254.84
+Table 1: FID score obtained by PhysGaussian [24] and GASP for static scenes and renders on images taken across simulations.
+For both methods, we use the same physics engine and trained GS model.
+In the case of GASP, we treated each triangle vertex as a Gaussian and used their positions across the simulation to calculate the actual parameters of each Gaussian.
+Method Vasedeck Fox Teddybear Ficus Bottle
+PhysGaussian 53.82 110.37 618.07 89.29 46.42
+GASP (our) 74.55 168.81 614.55 186.44 87.03
+GASP (our with Core) 17.94 64.69 351.18 19.93 18.80
+Table 2: Comparison of time in seconds required to finish the simulation for various models between PhysGaussian and GASP.
+Both methods use the same GS models and physics engine.
+For GASP, we also show the speed for simulations where only Core Gaussians were used during simulation.
+Figure 13: Comparison of renders obtained by GASP and PhysGaussian on ficus and fox simulations.
+For both methods, the same trained model was used, and the simulations were conducted using the physics engine used in PhysGaussian.
+For GASP we used vertices of triangle soup as positions used during simulation, after each step their position was used to calculate the parameters of Gaussians represented by each triangle.
+$\alpha$ 1.0 2.0 3.0 4.0 5.0 $\infty$
+FID 156.99 162.91 168.21 168.73 168.59 188.21
+Table 3: Ablation study of the proposed strategy with varying $\alpha$ values.
+The FID score generally increases with larger $\alpha$, highlighting the effectiveness of the method.
+$\infty$ indicates the strategy was not applied.
+We conduct an ablation study to assess the impact of the proposed Gaussian Control Rule.
+The results, summarized in Table 3, report the value of $\alpha$ used in Equation 12 (first row) and the corresponding FID score on the simulation shown in Figure 7 (second row).
+The findings indicate that the strategy yields consistent improvements.
+Furthermore, the FID score generally increases with larger $\alpha$ values, with only a slight drop observed between $\alpha=4.0$ and $\alpha=5.0$.
 5 Conclusion
-In this paper, we introduce GASP, a model that integrates high-quality rendering with accurate physical simulation.
-By leveraging a GaMeS-based representation of objects, our approach enables the simulation of basic phenomena, such as crushing, and more complex scenarios, including cracking and rainfall.
-This demonstrates the model’s versatility and effectiveness in handling a wide range of physical interactions within simulated environments.
-Experimental section, shows that GASP work in many difference scenario and produce high quality simulations.
+In this paper, we introduce GASP, a pipeline that integrates high-quality rendering with accurate physics simulation.
+Using a GaMeS-based representation of objects, our approach enables the simulation of basic phenomena, such as crushing, and more complex scenarios, including cracking.
+This demonstrates the model’s versatility and effectiveness in handling a wide range of physics interactions within simulated environments.
+Conducted experiments show that GASP works in many different scenarios and produces high-quality simulations.
+We observe that the additional control rules we proposed for hierarchical simulations significantly accelerate the simulation process.
+This approach resembles the particle sampler commonly utilized in many physics engines.
 Limitations
-Our model produces artifacts when the simulation changes shapes essentially.
+Our model may produce artifacts when the simulation causes drastic changes in the shape of the original object.
+Moreover, the GASP pipeline does not account for variations in lighting interaction with the modified surface.
+Since our method is built upon the GaMeS representation, it also inherits its limitations.
+Flat Gaussians may require a large number of components to accurately capture highly curved or anisotropic regions, while the mesh-free triangle soup lacks explicit connectivity.
+On the other hand, in mesh-based settings, the representation quality is constrained by the resolution of the mesh.
+6 Acknowledgements
+The project “Effective rendering of 3D objects using Gaussian Splatting in an Augmented Reality environment” (FENG.02.02-IP.05-0114/23) is carried out within the First Team programme of the Foundation for Polish Science co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG).
+The work of P.
+Spurek was supported by the National Science Centre (Poland), Grant No.
+2021/43/B/ST6/01456.